Accounting for Bonds

On May 1, 2020, Merrill Corp issued $1 million ten-year bonds. The bonds pay 10% interest semi-annually on October 31 and April 30. The bonds were issued to yield 12%. Merrill Corp has a fiscal year that ends December 31 each year. Fields Corporation uses the effective-interest method to calculate its interest expense each period.

Required:

  1. Compute the issue price of the bonds and the journal entry at issuance.
  2. Prepare the bond amortization table through April 30, 2021.
  3. Prepare the journal entry or entries associated with this bond on December 31, 2020 and April 30, 2021.
  4. Assume on May 1, 2021, Merill Corp decided to early retire the bonds by paying at 97 (97% of the face amount). Prepare the journal entry for bond early retirement.

PV of ordinary annuity

Period

4%

5%

6%

8%

10%

12%

10

8.11090

7.72173

7.36009

6.71008

6.14457

5.65002

20

13.59033

12.46221

11.46992

9.81815

8.51356

4.86958

PV of $1

Period

4%

5%

6%

8%

10%

12%

10

0.67556

0.61391

0.55839

0.46319

0.38554

0.32197

20

0.45639

0.37689

0.31180

0.21455

0.14864

0.10367
Answer #1

Solution

Merrill Corp

Computation of the issue price of bonds and the journal entry –

Issue price of bonds = present value of bonds + present value of interest payments

Present value of bond = face value x (P/F, r, N)

Effective rate of interest, r = 12% x 6/12 = 6%

Period, N = 10 years x 2 semi-annual periods = 20

Face value = $1,000,000

Present value of bonds = 1,000,000 x (P/F, 6%, 20)

= 1,000,000 x 0.31180 = $311,800

Present value of interest payments –

Semi-annual interest payment = 1,000,000 x 10% x 6/12 = $50,000

Present value of interest payments = 50,000 x (P/A, 6%, 20)

= 50,000 x 11.46992 = $573,496

Issue price of bonds = $311,800 + 573,496 = $885,296

Since, the issue price is less than the face value, the bond is said to have been issued at discount.

Discount on Bonds Payable = 1,000,000 – 885,296 = $114,704

Journal entry –

Account Titles and Explanation Debit Credit Date May 1, 2020 $885,296 Cash Discount on Bonds Payable Bonds Payable (To record

Interest Expense Discount Amortization Carrying Value of Bond $885,296 Bond Amortization Table Cash Date Payment May 1, 2020

Computations –

Interest expense = carrying value of bond x effective market rate

Cash payment = face value x interest rate x semi-annual period

Discount Amortization = interest expense – cash payment

Carrying value of bond = beg. Balance + discount amortization

Oct 31, 2020 –

Interest expense = 885,296 x 6% = 53,118

Cash payment = 1,000,000 x 10% x 6/12 = $50,000

Discount Amortization = 53,118 – 50,000 = $3,118

Carrying value of bond = 884,296 + 3,118 = $888,414

April 30, 2021 –

Interest expense = 888,414 x 6% = $53,305

Cash payment = $50,000

Discount amortization = 53,305 – 50,000 = $3,305

Carrying Value of Bond = 888,414 + 3,305 = $891,719

Journal entries on Dec 31, 2020 and April 30, 2021 –

Account Titles and Explanation Debit Credit Date Dec 31, 2020 Interest Expense $17,768 Discount on Bonds Payable $1,101 Inter

Computations –

The next interest payment date is on April 30. However, on Dec 31, the company recognizes the interest accrued as follows,

Dec 31, 2020

Interest expense = 888,414 x 6% x 2/6 = 17,768

Interest payable = 1,000,000 x 10% x 6/12 x 2/6 = 16,667

Discount on bonds payable = 17,768 – 16,667 = 1,101

April 30, 2021

Interest expense = 888,414 x 6% x 4/6 = 35,537

Since, the accrued interest is paid, cash $50,000

Discount on bonds payable = 35,537 + 16,667 – 50,000 = 2,204

Retirement of bond on May 1, 2021 –

Journal entry for bond retirement –

Account Titles and Explanation Debit Credit Date May 1, 2021 $891,719 $186,562 Bonds Payable Loss on Retirement of Bonds Disc

Computations -

Carrying value of bond = 891,719

Cash payment = 97% x 1,000,000 = 970,000

Unamortized discount = 114,704 – (3,118 +3,305) = 108,281

Loss on retirement = 970,000 + 108,281 – 891,719 = 186,562